₦airaland Forum

Welcome, Guest: Register • Login • With Google • Trending • Recent • New

Stats: 3,344,286 members, 8,511,065 topics. Date: Saturday, 31 October 2026 at 05:41 PM

Toggle theme

Probability Help - Education - Nairaland

Nairaland Forum › Nairaland General › Education › Probability Help (677 Views)

1 Reply (Go Down)

Probability Help by sniperwolf(op): 4:34pm On Oct 29, 2011
Please I need help with this probability question:

ONE WHITE DIE AND ONE BLACK DIE ARE TOSSED. FIND THE PROBABILITY THAT THE WHITE DIE SHOWS A NUMBER SMALLER THAN 5 OR THAT THE SUM OF BOTH DICE IS GREATER THAN 8.

N/B: Please I need workings
Re: Probability Help by Dstar4u(m): 6:31pm On Oct 29, 2011
4/6 or 1/12. Thisz d wokin: prob dat d white die tossed is smaller dan 5 is 1, 2, 3 & 4, it means d no smaller dan 5 is 4 dere4 d prob is 4/6. Or d prob of d sum is equal to 8
Re: Probability Help by sniperwolf(op): 8:42am On Oct 30, 2011
What's that? The event is Non-Mutually Exclusive. I don't think you know what you did
Re: Probability Help by 4ortyse7en(m): 1:02pm On Oct 30, 2011
The probability that the white die shows a number less than 5 is = 4/6 = 2/3.
For the sum of the numbers shown on both dice to be greater than 8, the possible combinations are (3,6) (4,5) (4,6) (5,4) (5,5) (5,6) (6,3) (6,4) (6,5) (6,6). Since there are 10 outcomes out of a total possible 36, thus the probability is 10/36= 5/18.
The probability that either of the two events will occur is = 2/3 + 5/18 = 17/18.
Re: Probability Help by 4ortyse7en(m): 1:05pm On Oct 30, 2011
The probability that the white die shows a number less than 5 is = 4/6 = 2/3.
For the sum of the numbers shown on both dice to be greater than 8, the possible combinations are (3,6) (4,5) (4,6) (5,4) (5,5) (5,6) (6,3) (6,4) (6,5) (6,6). Since there are 10 combinations out of a total possible 36, thus the probability is 10/36= 5/18.
The probability that either of the two events will occur is = 2/3 + 5/18 = 17/18.
Re: Probability Help by 4ortyse7en(m): 2:02pm On Oct 30, 2011
After looking at the question more thoughtfully again, I discovered that since the events are non mutually exclusive, the probability that either of the events will occur = P(A) + P(B) - P(A & B).
Where P(A)= 2/3, P(B)= 5/18, P(A & B) = 3/36 because in 3 combinations of the 10 combinations I cited earlier, the white die shows a number that's below 5. This is what makes it non mutually exclusive.
Therefore the correct probability is = 2/3 + 5/18 - 3/36 = 31/36.
Hope this helps.
1 Reply

Mathematicians In The House, Please I Need Solution To This Probability Problem • Statistics/probability Questions. • Pls Can U Solve This Probability • 2 • 3 • 4

Pls, Advise On How To Pursue This Injustice • Government, ASUU Negotiationto End Soon, Says Labourministerdecember 29, 2011 04:27pmprint Email Sth • Students Of Sociology/Anthropology